# ViMultiChoice: Toward a Method That Gives Explanation for Multiple-Choice Reading Comprehension in Vietnamese

### v1

@@ Abstract
Multiple-choice Reading Comprehension (MCRC) models aim to select the correct answer from a set of candidate options for a given question. However, they typically lack the ability to explain the reasoning behind their choices. In this paper, we introduce a novel Vietnamese dataset designed to train and evaluate MCRC models with explanation generation capabilities. Furthermore, we propose ViMultiChoice, a new method specifically designed for modeling Vietnamese reading comprehension that jointly predicts the correct answer and generates a corresponding explanation. Experimental results demonstrate that ViMultiChoice outperforms existing MCRC baselines, achieving state-of-the-art (SotA) performance on both the ViMMRC 2.0 benchmark and the newly introduced dataset. Additionally, we show that jointly training option decision and explanation generation leads to significant improvements in multiple-choice accuracy.
@@ 1 Introduction
Multiple-choice Reading Comprehension (MCRC) plays an important role in a wide range of applications, including virtual assistants and conversational agents. Recent advances in Information Retrieval (IR) Zeng et al. (2024) and Large Language Models (LLMs) et al. (2024) have opened up new directions for research and development in this area.
While MCRC research has been extensively studied for high-resource languages such as English and Chinese, work on low-resource languages, including Vietnamese, remains limited. Existing studies on Machine Reading Comprehension (MRC) in general, and MCRC in particular, for Vietnamese have primarily focused on introducing large-scale benchmarks across different domains, such as healthcare news Nguyen et al. (2020b); Luu et al. (2021) and education Luu et al. (2023). In addition, most proposed Vietnamese MCRC models rely on complex architectures that integrate multiple components, including Natural Language Inference (NLI) Camburu et al. (2018); Bowman et al. (2015) and multi-hop question answering Yang et al. (2018). These approaches are computationally expensive and often overlook the linguistic characteristics specific to Vietnamese. Moreover, MCRC methods that are capable of providing explanations for their predictions are more desirable, as they offer greater transparency and are more suitable for real-world applications.
@@ 2.1 Datasets
Recent advances in Machine Reading Comprehension (MRC), particularly in the multiple-choice setting (MC-MRC), have been driven by the release of large-scale benchmark datasets designed to evaluate increasingly complex reasoning abilities of language models. Early datasets such as MCTest Richardson et al. (2013) focused on basic reading comprehension but were limited in scale and reasoning depth, while RACE Lai et al. (2017) marked a significant step forward by introducing longer and more complex passages from high school examinations. However, subsequent studies revealed that many RACE questions could still be solved using shallow heuristics or by exploiting statistical answer-position biases. To address these limitations, newer benchmarks have been developed to target higher-level reasoning, including formal logical reasoning in ReClor Yu et al. (2020) or commonsense reasoning in CommonsenseQA Talmor et al. (2019).
In parallel, research on MC-MRC for Vietnamese has gained increasing attention, although the data ecosystem remains relatively limited compared to high-resource languages. Early Vietnamese MRC efforts such as UIT-ViQuAD Van Nguyen et al. (2020) focused on extractive question answering and did not capture the challenges of multiple-choice reasoning. Subsequent datasets, including ViMMRC 2.0 Luu et al. (2023), addressed this gap in educational contexts by introducing literary texts with higher-level reasoning requirements and semantically similar distractors. Despite these advances, existing Vietnamese datasets still face limitations in terms of reasoning depth, domain coverage, and explanation-based evaluation, highlighting the need for more comprehensive and explanation available MC-MRC benchmarks.
@@ 2.2 Methods
Multiple-choice Machine Reading Comprehension (MC-MRC) poses unique challenges compared to extractive or generative question answering, as it requires models to compare semantically similar options, perform multi-step reasoning, and eliminate distractors. Early approaches to MC-MRC largely inherited traditional MRC architectures based on sequential text representations, employing RNN-based models such as LSTMs or GRUs to independently encode the context, question, and each answer option. Representative methods from this stage include Memory Network–based models Tang et al. (2019), which treat textual units as memory slots and select answers by measuring relevance between question–option pairs and contextual memory, as well as CNN-based approaches Chaturvedi et al. (2018) that capture local semantic patterns. However, these methods generally lack explicit mechanisms for modeling fine-grained interactions among answer choices and struggle with long-range dependencies and complex reasoning.
To overcome these limitations, subsequent research shifted toward interaction- and matching-based methods that explicitly model semantic relationships between the context, question, and answer options. Co-matching frameworks such as Co-Match Wang et al. (2018) and MMN Tang et al. (2019) introduced joint and multi-level matching mechanisms to capture both local and global semantic alignments, while hierarchical and multi-paragraph models such as HAF Zhu et al. (2018) and DynSAN Zhuang and Wang (2019) addressed the challenges posed by long or multi-document contexts through hierarchical and dynamic attention mechanisms. In parallel, option-centric approaches including OCN Ran et al. (2019) and ElimiNet Parikh et al. (2018) focused on direct comparison and elimination of answer choices to better distinguish closely related distractors, mimicking human test-taking strategies. More recent pre-Transformer methods further enhanced matching architectures by modeling higher-order interactions, such as DCMN+ Zhang et al. (2020) and CTM Yao et al. (2021), which introduced strengthened co-matching and triple-matching mechanisms among context, question, and options. Overall, the evolution of MC-MRC methods reflects a progression from independent representation learning toward increasingly sophisticated interaction modeling and reasoning strategies.
@@ 3.1 Dataset Construction
Multiple-choice Questions Collection
@@
Answers and Explanation Annotation
The correct answers and their corresponding explanations were annotated by teachers from Nguyen Huu Huan high school. In total, 7 high school teachers participated in the annotation process, providing answers and explanations for all questions. For each subject, between two and three teachers served as annotators. All annotators have at least 10 years of teaching experience at the high school level and a minimum of 5 years of experience in preparing students for university entrance examinations. Within each subject, annotators cross-validated one another’s annotations to assess inter-annotator consistency. At the conclusion of the annotation process, the Cohen’s kappa inter-annotator agreement Cohen (1960) reached 96%, indicating a very high level of agreement among teachers across the four subjects in the ViRCSoSciD dataset.
Textbooks Collection
@@ 3.2 Overall Statistics
Overall, our dataset contains 12,819 multiple-choice questions. As shown in Figure 1, approximately half of the questions are designed for Grade 12 students, while Grade 10 accounts for the smallest proportion. Questions targeting Grade 12 students are generally more complex, as they often require comprehensive knowledge accumulated across all three high school grades to identify relevant information and determine the correct answer. These questions are primarily intended to support preparation for the Vietnamese university entrance examination.
@@ 3.3 On the Bias Choice Problem
A significant challenge in constructing multiple-choice question answering datasets is the presence of choice bias Mihaylov et al. (2018), where models tend to rely on the prior distribution of options (A, B, C, or D) rather than truly understanding the content of the question. To address this, we examined the frequency distribution of correct options and applied a de-biasing strategy by shuffling the order of choices. According to Figure 2, the distribution of four options is mostly uniform. There is an exception on Civic Education where the option A has the highest percentage. However, the distribution of four option is still uniform in general (Figure 1(a)), which directly limit the effect of the bias choice problem.
@@ 4.1 Overall Architecture
Our proposed ViMultiChoice (Multiple-Choice Reading Comprehension in Vietnamese) is developed based on the OCN Ran et al. (2019) method with the advancement in Vietnamese language modeling and explanation generation. In particular, the ViMultiChoice has three main components: Vietnamese Text Encoder, Option Inference and Explanation Generator (Figure 3).
@@ 4.2 Vietnamese Text Encoder
Vietnamese Text Encoder aims to encode and represent Vietnamese texts in questions, options, and contexts for answer inference and explanation generation. This module consists of two components: (1) a pretrained language model for Vietnamese and (2) the ViWordFormer module.
@@
We inspired by the recent study Hieu Nguyen et al. (2025) to have a module that describes the linguistic features of Vietnamese phrasal lexemes for better understanding and representation modeling in this language. In particular, Vietnamese has compound words having two syllables separated by spaces in writing. To this end, Nguyen et al. (2009) claimed that word segmentation is a linguistic feature of Vietnamese. However Hieu Nguyen et al. (2025) argue that such compound words are actually phrases in Vietnamese, hence word segmentation is a particular way of modeling Vietnamese phrases. Following this study, we propose the ViWordFormer module which describes phrasal lexemes in Vietnamese texts.
Given the embedded vectors of the input text $w=\{w_{1},w_{2},...,w_{n_{w}}\}$ as $f_{w}=[f^{w}_{1},f^{w}_{2},...,f^{w}_{n_{w}}]^{T}\in\mathbb{R}^{n_{w}\times d}$, ViWordFormer first has the multihead self-attention module of Transformer Vaswani et al. (2017) for determining the Attention Score matrix $A$ of every token in $f_{w}$:
@@
where $P_{k}$ is the probability of $w_{k}$ and $w_{k+1}$ are in the same phrase or they form a compound word Nguyen et al. (2009). This term is calculated as the geometric average of two terms $pr_{k,k+1}$ and $pr_{k+1,k}$:
@@ 4.3 Option Inference Module
Given the features $f_{Q},f_{O_{k}},f_{P}$ from the Vietnamese Word Encoder, we develop the Option Inference Module which is mostly based on the OCN Ran et al. (2019) method for reasoning and deciding options. This module contains four Option Comparison module. Each of them determines the feature vector $f_{k}$ representing for each option $O_{k}$ after considering the information from other options and the context. The description of this module is provided in detailed in the Appendix section due to the limitation of page allowance.
@@ 5.1 Datasets and Metrics
Datasets.
We evaluated our proposed method against other baselines in two MCRC corpora in Vietnamse: ViMMRC 2.0 Luu et al. (2023) and our ViRCSoSciD dataset. The ViMMRC 2.0 dataset contains 5,273 multiple-choice questions in literature for students from primary school to high school in Vietnam. On the ViMMRC dataset we evaluated the reading comprehension by determining the correct option while on the ViRCSoSciD we evaluated both reading comprehension and explanation ability.
Metrics.
We followed previous studies Luu et al. (2023) to have F1-macro and Accuracy for evaluating the baselines and our proposed method in multiple-choice option decision. For explanation generation task, we followed studies in text summarization to deploy BLEU-4 Papineni et al. (2002) and ROUGE-L Lin (2004) to measure how much the explanation given by baselines and our ViMultiChoice is closed to the ground truth explanation.
@@ 5.2 Baselines
On our ViRCSoSciD dataset, we evaluated two classes of baselines: Large Language Models (LLMs) and the standard neural-based methods for MCRC. In particular, we evaluated open-source LLMs which have ability of reading and understanding Vietnamese: Germini-Flash Team (2025), Mistral-large et al. (2023), QWen2 et al. (2024), Vicuna Zheng et al. (2023), URA-LlaMA Truong et al. (2024), Vina-LlMA Nguyen et al. (2023), and Flan-T5 et al. (2022). These LLMs are zero-shot prompted on the ViRCSoSciD dataset to perform two tasks: multiple-choice reading comprehension and explanation.
@@
  • Bidirectional Semantic Matching: DCMN+ Zhang et al. (2020), CNN for MCRC Chaturvedi et al. (2018), HAF Zhu et al. (2018), MMN Tang et al. (2019), and DynSAN Zhuang and Wang (2019).
  • Three-Directional Matching: Co-Match Wang et al. (2018) and CTM Yao et al. (2021).
  • Option Comparison and Elimination: OCN Ran et al. (2019) and ElimiNet Parikh et al. (2018).

### v2

# arXiv:2602.09961v2[cs.CL] 27 Feb 2026

@@ Abstract
Multiple-choice Reading Comprehension (MCRC) models aim to select the correct answer from a set of candidate options for a given question. However, they typically lack the ability to explain the reasoning behind their choices. In this paper, we introduce a novel Vietnamese dataset designed to train and evaluate MCRC models with explanation generation capabilities. Furthermore, we propose ViMultiChoice, a new method specifically designed for modeling Vietnamese reading comprehension that jointly predicts the correct answer and generates a corresponding explanation. Experimental results demonstrate that ViMultiChoice outperforms existing MCRC baselines, achieving state-of-the-art (SotA) performance on both the ViMMRC 2.0 benchmark and the newly introduced dataset. Additionally, we show that jointly training option decision and explanation generation leads to significant improvements in multiple-choice accuracy. To support reproducibility, the code and dataset will be publicly released upon acceptance of this paper.
keywords:
Multiple-choice , Question Answering , Explanation , Vietnamese , Transformer
journal: International Journal of Artificial Intelligence in Education
[…]
[1]organization=Faculty of Information Science and Engineering
[…]
[2]organization=Faculty of Computer Science
[…]
[3]organization=University of Information Technology, Vietnam National University
[…]
[4]organization=Vietnam National University, city=Ho Chi Minh city, country=Viet Nam
@@ 1 Introduction
Multiple-choice Reading Comprehension (MCRC) plays an important role in a wide range of applications, including virtual assistants and conversational agents. Recent advances in Information Retrieval (IR) [43] and Large Language Models (LLMs) [38] have opened up new directions for research and development in this area.
While MCRC research has been extensively studied for high-resource languages such as English and Chinese, work on low-resource languages, including Vietnamese, remains limited. Existing studies on Machine Reading Comprehension (MRC) in general, and MCRC in particular, for Vietnamese have primarily focused on introducing large-scale benchmarks across different domains, such as healthcare news [19, 14] and education [15]. In addition, most proposed Vietnamese MCRC models rely on complex architectures that integrate multiple components, including Natural Language Inference (NLI) [2, 1] and multi-hop question answering [39, 5]. These approaches are computationally expensive and often overlook the linguistic characteristics specific to Vietnamese. Moreover, MCRC methods that are capable of providing explanations for their predictions are more desirable, as they offer greater transparency and are more suitable for real-world applications.
@@
Through these contributions, we aim to provide a foundational framework for future research on MCRC modeling for Vietnamese and other low-resource languages.
@@ 2.1 Datasets
Recent advances in Machine Reading Comprehension (MRC), particularly in the multiple-choice setting (MC-MRC), have been driven by the release of large-scale benchmark datasets designed to evaluate increasingly complex reasoning abilities of language models. Early datasets such as MCTest [28] focused on basic reading comprehension but were limited in scale and reasoning depth, while RACE [11] marked a significant step forward by introducing longer and more complex passages from high school examinations. However, subsequent studies revealed that many RACE questions could still be solved using shallow heuristics or by exploiting statistical answer-position biases. To address these limitations, newer benchmarks have been developed to target higher-level reasoning, including formal logical reasoning in ReClor [41] or commonsense reasoning in CommonsenseQA [30], and multi-step scientific reasoning in OpenBookQA [17]. More recent datasets, such as MMMU [42] and MathVista [13], further extend MC-MRC to expert-level academic and mathematical domains, reflecting a clear shift from surface-level text understanding toward structured, multi-step reasoning.
In parallel, research on MC-MRC for Vietnamese has gained increasing attention, although the data ecosystem remains relatively limited compared to high-resource languages. Early Vietnamese MRC efforts such as UIT-ViQuAD [34] focused on extractive question answering and did not capture the challenges of multiple-choice reasoning. Subsequent datasets, including ViMMRC 2.0 [15], addressed this gap in educational contexts by introducing literary texts with higher-level reasoning requirements and semantically similar distractors. Despite these advances, existing Vietnamese datasets still face limitations in terms of reasoning depth, domain coverage, and explanation-based evaluation, highlighting the need for more comprehensive and explanation available MC-MRC benchmarks.
@@ 2.2 Methods
Multiple-choice Machine Reading Comprehension (MC-MRC) poses unique challenges compared to extractive or generative question answering, as it requires models to compare semantically similar options, perform multi-step reasoning, and eliminate distractors. Early approaches to MC-MRC largely inherited traditional MRC architectures based on sequential text representations, employing RNN-based models such as LSTMs or GRUs to independently encode the context, question, and each answer option. Representative methods from this stage include Memory Network–based models [31], which treat textual units as memory slots and select answers by measuring relevance between question–option pairs and contextual memory, as well as CNN-based approaches [3] that capture local semantic patterns. However, these methods generally lack explicit mechanisms for modeling fine-grained interactions among answer choices and struggle with long-range dependencies and complex reasoning.
To overcome these limitations, subsequent research shifted toward interaction- and matching-based methods that explicitly model semantic relationships between the context, question, and answer options. Co-matching frameworks such as Co-Match [37] and MMN [31] introduced joint and multi-level matching mechanisms to capture both local and global semantic alignments, while hierarchical and multi-paragraph models such as HAF [46] and DynSAN [47] addressed the challenges posed by long or multi-document contexts through hierarchical and dynamic attention mechanisms. In parallel, option-centric approaches including OCN [26] and ElimiNet [24] focused on direct comparison and elimination of answer choices to better distinguish closely related distractors, mimicking human test-taking strategies. More recent pre-Transformer methods further enhanced matching architectures by modeling higher-order interactions, such as DCMN+ [44] and CTM [40], which introduced strengthened co-matching and triple-matching mechanisms among context, question, and options. Overall, the evolution of MC-MRC methods reflects a progression from independent representation learning toward increasingly sophisticated interaction modeling and reasoning strategies.
@@ 3.1 Dataset Construction
3.1.1 Multiple-choice Questions Collection
@@
3.1.2 Answers and Explanation Annotation
The correct answers and their corresponding explanations were annotated by teachers from Nguyen Huu Huan high school. In total, 7 high school teachers participated in the annotation process, providing answers and explanations for all questions. For each subject, between two and three teachers served as annotators. All annotators have at least 10 years of teaching experience at the high school level and a minimum of 5 years of experience in preparing students for university entrance examinations. Within each subject, annotators cross-validated one another’s annotations to assess inter-annotator consistency. At the conclusion of the annotation process, the Cohen’s kappa inter-annotator agreement [7] reached 96%, indicating a very high level of agreement among teachers across the four subjects in the ViRCSoSciD dataset.
3.1.3 Textbooks Collection
@@ 3.2 Overall Statistics
Figure 2: Some samples from the ViRCSoSciD dataset.
Overall, our dataset contains 12,819 multiple-choice questions. As shown in Figure 1, approximately half of the questions are designed for Grade 12 students, while Grade 10 accounts for the smallest proportion. Questions targeting Grade 12 students are generally more complex, as they often require comprehensive knowledge accumulated across all three high school grades to identify relevant information and determine the correct answer. These questions are primarily intended to support preparation for the Vietnamese university entrance examination (Figure 2).
@@ 3.3 On the Bias Choice Problem
A significant challenge in constructing multiple-choice question answering datasets is the presence of choice bias [16], where models tend to rely on the prior distribution of options (A, B, C, or D) rather than truly understanding the content of the question. This bias undermines the model’s ability to reason correctly and can lead to inaccuracies in performance evaluations. To address this, we examined the frequency distribution of correct options and applied a de-biasing strategy by shuffling the order of choices. This process reduces the impact of label position shortcuts and creates a more balanced dataset, ensuring that model performance is based on true knowledge and reasoning rather than statistical patterns.
According to Figure 3, the distribution of four options is mostly uniform. There is an exception on Civic Education where the option A has the highest percentage. However, the distribution of four option is still uniform in general (Figure 1(a)), which directly limit the effect of the bias choice problem.
@@ 4.1 Overall Architecture
Our proposed ViMultiChoice (Multiple-Choice Reading Comprehension in Vietnamese) is developed based on the OCN [26] method with the advancement in Vietnamese language modeling and explanation generation. In particular, the ViMultiChoice has three main components: Vietnamese Text Encoder, Option Inference and Explanation Generator (Figure 4).
4.2 Context Retrieval Module
We applied an information retrieval pipeline to retrieve supporting context for multiple-choice questions. The process begins by filtering knowledge based on the question’s subject, followed by splitting the relevant content into sentence-level units. To enhance retrieval precision, we construct queries by averaging the representations of the question and its corresponding answer options. This approach ensures that the retrieved context captures the collective semantics of the entire multiple-choice item, rather than being limited to the question’s isolated information.
In this study, we examined the effectiveness of SotA text retrievers to collect relevant context from text books for the questions of the ViRCSoSciD corpus. In particular, we conducted experiments for BM25 [29], MiniLML12 [27], MultiE5 [36], VietBGE [20], and VietBGEMultiE5 which is the combination of VietBGE and MultiE5 pretrained embeddings. Experimental results show that VietbGE is the most effective one. From that on, MCRC methods on the ViRCSoSciD were trained with context retrieved by VietBGE from the Vietnamese standard textbooks.
@@ 4.3 Vietnamese Text Encoder
Figure 5: The architecture of Vietnamese Text Encoder. MSA stands for Multihead Self-Attention module
Vietnamese Text Encoder aims to encode and represent Vietnamese texts in questions, options, and contexts for answer inference and explanation generation. This module consists of two components: (1) a pretrained language model for Vietnamese and (2) the ViWordFormer module (Figure 5).
@@
We inspired by the recent study [8] to have a module that describes the linguistic features of Vietnamese phrasal lexemes for better understanding and representation modeling in this language. In particular, Vietnamese has compound words having two syllables separated by spaces in writing. To this end, [21] claimed that word segmentation is a linguistic feature of Vietnamese. However [8] argue that such compound words are actually phrases in Vietnamese, hence word segmentation is a particular way of modeling Vietnamese phrases. Following this study, we propose the ViWordFormer module which describes phrasal lexemes in Vietnamese texts.
Given the embedded vectors of the input text $w=\{w_{1},w_{2},...,w_{n_{w}}\}$ as $f_{w}=[f^{w}_{1},f^{w}_{2},...,f^{w}_{n_{w}}]^{T}\in\mathbb{R}^{n_{w}\times d}$, ViWordFormer first has the multihead self-attention module of Transformer [35] for determining the Attention Score matrix $A$ of every token in $f_{w}$:
@@
where $P_{k}$ is the probability of $w_{k}$ and $w_{k+1}$ are in the same phrase or they form a compound word [21]. This term is calculated as the geometric average of two terms $pr_{k,k+1}$ and $pr_{k+1,k}$:
@@ 4.4 Option Inference Module
Figure 6: The Option Inference module. OptCom stands for Option Comparison.
Given the features $f_{Q},f_{O_{k}},f_{P}$ from the Vietnamese Word Encoder, we develop the Option Inference Module which is mostly based on the OCN [26] method for reasoning and deciding options. This module contains four Option Comparison module. Each of them determines the feature vector $f_{k}$ representing for each option $O_{k}$ after considering the information from other options and the context (Figure 6).
$f_{O_{k}}^{Q}=[f_{Q};f_{O_{k}}]\in\mathbb{R}^{n_{Qk}\times d},k\in\{1,2,3,4\}$
where $[;]$ denotes the concatenation operator along the last dimension and $n_{Qk}=n_{Q}+n_{k}$. First, we compare any option $O_{k}$ with another option $O_{l}(l\neq k)$ via the trilinear attention mechanism [26]:
$f_{O_{k,l}}^{Q}=Attn(f_{O_{k}}^{Q},f_{O_{l}}^{Q})f_{O_{l}}^{Q}$ (7)
where $W_{C}\in\mathbb{R}^{4d\times d}$ and $b_{C}\in\mathbb{R}^{d}$.
$f^{O}_{k}=g_{k}\odot f_{O_{k}}^{Q}+(1-g_{k})\odot\bar{f}^{Q}_{O_{k}}$ (12)
$A^{O}_{k}=Attn(f_{k}^{O},f_{P})\in\mathbb{R}^{n_{Qk}\times n_{P}}$ (14)
$A^{P}_{k}=Attn(f_{P},f_{k}^{O})\in\mathbb{R}^{n_{P}\times n_{Qk}}$ (15)
$\hat{f}^{O}_{k}=A^{O}_{k}[f_{P};A^{P}_{k}f^{O}_{k}]\in\mathbb{R}^{n_{Qk}\times 2d}$ (16)
where
and
@@ 5.1 Datasets and Metrics
5.1.1 Datasets
We evaluated our proposed method against other baselines in two MCRC corpora in Vietnamese: ViMMRC 2.0 [15] and our ViRCSoSciD dataset. The ViMMRC 2.0 dataset contains 5,273 multiple-choice questions in literature for students from primary school to high school in Vietnam. On the ViMMRC dataset we evaluated the reading comprehension by determining the correct option while on the ViRCSoSciD we evaluated both reading comprehension and explanation ability.
5.1.2 Metrics
We followed previous studies [15] to have F1-macro and Accuracy for evaluating the baselines and our proposed method in multiple-choice option decision. For explanation generation task, we followed studies in text summarization to deploy BLEU-4 [23] and ROUGE-L [12] to measure how much the explanation given by baselines and our ViMultiChoice is closed to the ground truth explanation.
@@ 5.2 Baselines
On our ViRCSoSciD dataset, we evaluated two classes of baselines: Large Language Models (LLMs) and the standard neural-based methods for MCRC. In particular, we evaluated open-source LLMs which have ability of reading and understanding Vietnamese: Gemini-Flash [32], Mistral-large [9], QWen2 [38], Vicuna [45], URA-LlaMA [33], Vina-LlaMA [22], and Flan-T5 [6]. These LLMs are zero-shot prompted on the ViRCSoSciD dataset to perform two tasks: multiple-choice reading comprehension and explanation.
@@
  • 1.
    Bidirectional Semantic Matching: DCMN+ [44], CNN for MCRC [3], HAF [46], MMN [31], and DynSAN [47].
  • 2.
    Three-Directional Matching: Co-Match [37] and CTM [40].
  • 3.
    Option Comparison and Elimination: OCN [26] and ElimiNet [24].